OpenStudy (mpj4):

how do you even solve this? 1) y = 2^x ; y = x + 1 and 2) y = e^-x ; y = x^2

OpenStudy (fibonaccichick666):

I guess my question is, what do you mean by solve?

OpenStudy (mpj4):

(x,y) coordinates

OpenStudy (fibonaccichick666):

well, if you just need some, just pick a few x values and plug them in where you see an x. x=0,-1,1 are usually good choices

OpenStudy (mpj4):

ah, then that solves 1, I guess there was no proper way to solve 1. How about 2?

OpenStudy (fibonaccichick666):

same thing?

OpenStudy (fibonaccichick666):

do you mean solve the system of equations ie where they are equal?

OpenStudy (mpj4):

Yeah.

OpenStudy (fibonaccichick666):

that's a useful bit of info

OpenStudy (mpj4):

Sorry, just found it in the book.

OpenStudy (fibonaccichick666):

so set \[e^{-x}=x^2\]

OpenStudy (fibonaccichick666):

(y=y)<-- that way you can find the answer, so then, solve for x or use a quick observation

OpenStudy (fibonaccichick666):

what is the biggest e^-x can ever be?

OpenStudy (mpj4):

0?

OpenStudy (fibonaccichick666):

2^0 doesn't equal zero does it?

OpenStudy (fibonaccichick666):

and asked for the biggest it could be

OpenStudy (fibonaccichick666):

also, another way is to just graph them and take a look at the point of intersection

OpenStudy (mpj4):

It's 1, sorry, it was either 1 or 0 when I looked at it. Didn't think it through xD

OpenStudy (fibonaccichick666):

*biggest it can be provided x is positive, **

OpenStudy (fibonaccichick666):

but yea, 1, does x^2 ever equal 1?

OpenStudy (mpj4):

yeah

OpenStudy (fibonaccichick666):

are the x values the same when it happens?

OpenStudy (mpj4):

no.

OpenStudy (fibonaccichick666):

ok so the algebraic way is messy, I recommend graphing it's easier. Here use this https://www.desmos.com/calculator

OpenStudy (mpj4):

I solved it with this but yeah, Algebraically would have been nice xD

OpenStudy (fibonaccichick666):

\[e^{-x}=x^2\] we'd have to use a log.... hmm -x=2lnx

OpenStudy (fibonaccichick666):

-x/lnx=2 ln e^-x/lnx=2 ln(e^{-x}-x)=2

OpenStudy (fibonaccichick666):

e^{-x}-x=e^2... hmm

OpenStudy (fibonaccichick666):

hmm.... or maybe since y=e^{-x} then \[x=-lny\] and \[x=\frac{+}{-}\sqrt{x}\]

OpenStudy (fibonaccichick666):

we get y=2lny then hmm

OpenStudy (fibonaccichick666):

still can't solve

OpenStudy (fibonaccichick666):

any ideas on an algebraic way guys? @ganeshie8 , @eliassaab , @hero, @ikram002p , @ wio

OpenStudy (mpj4):

You don't have to solve that second one anymore, unless you want to. I just saw it and wanted to know how.

OpenStudy (fibonaccichick666):

I do too... that's the thing... I can't figure it out

OpenStudy (anonymous):

\[ y=x+1; \quad y= 2^x\\ 2^x=x+1\\ x=0 \text { or } x=1 \]

OpenStudy (mpj4):

Oh, thanks.

ganeshie8 (ganeshie8):

familiar with lambert w function ?

ganeshie8 (ganeshie8):

@Kainui

OpenStudy (anonymous):

For the second one there no really elementary algebraic solution unless you use the productlog function. A numerical solution will give x= .7034

OpenStudy (kainui):

\[\LARGE e^{-x}=x^2 \]\[\LARGE 1 = x^2 e^x\]\[\LARGE \pm 1 = xe^{x/2} \]\[\LARGE \frac{\pm 1}{2}=\frac{x}{2}e^{x/2} \]\[\LARGE W(\frac{\pm1}{2})=\frac{x}{2}\]\[\LARGE 2 W(\frac{\pm 1}{2})=x\] So although this is one answer: http://www.wolframalpha.com/input/?i=2*W(1%2F2)&t=crmtb01 the other answer is a complex number since the smallest real value is -1/e which is close to -1/3: http://www.wolframalpha.com/input/?i=2*W%28-1%2F2%29

OpenStudy (mpj4):

Ugghh.. I haven't a clue on how all that happened @.@ Don't bother explaining, I've already wasted your time xD

OpenStudy (fibonaccichick666):

thanks guys, I was thoroughly curious!